OpenStudy (anonymous):

Find the x-intercept for this equation: 3x - 5y = 30

OpenStudy (dayakar):

to find x- intercept write y=0 in the equation

OpenStudy (dayakar):

3x -5*0 = 30

OpenStudy (dayakar):

3x= 30

OpenStudy (dayakar):

x= 30/3 x=10

OpenStudy (dayakar):

is it clear

OpenStudy (dayakar):

x-intercept = (10,0)

OpenStudy (anonymous):

@dayakar yes thank you